write the equation of the line passing through (4,4) and (4,2)
A) y=4
B)y=-2x
C) x=4
D) y=x+4

Answer by Nate(3500)

(4,4) and (4,2)
M = (y2 - y1)/(x2 - x1) = (4 - 2)/(4 - 4) = 2/0 the line is undefined, so the line is vertical
x+=+4
(C)